OpenStudy (anonymous):

Are the graphs of the lines in the pair parallel? Explain. y=2\3x-17 4x-6y=-6

OpenStudy (anonymous):

yes because both equations have 2/3x which makes them parallel

OpenStudy (anonymous):

Lines are parallel if their slopes are the same. Write each equation in the form y = mx + c. If m is the same the two lines are parallel. So write the equations using y=mx+b y=2/3x-17 is already in that form 4x-6y=-6 is not. To get that equation in the formula, subtract 4x from both sides so the -6y=-4x-6. Next divide by -6 so that y= -4/-6x+1. Youre not done yet though! :) simplify the sslope -4/-6 so that the final equation is y=2/3x+1 :) Therefore they have the same slopes

OpenStudy (mathmale):

"yes because both equations have 2/3x which makes them parallel": You are comparing SLOPES. The SLOPES of the two lines are the same: 2/3. Leave that "x" out.